OpenStudy (anonymous):

can someone help explain how to solve quadratic equations by completing the square? youtube videos aren't really helping me

jimthompson5910 (jim_thompson5910):

Do you have an example you want to work with?

OpenStudy (anonymous):

h(t)=-16t^2+576t

jimthompson5910 (jim_thompson5910):

Ok for this example, I'm going to use x in place of t. So instead of h(t)=-16t^2+576t, I'm going to use h(x)=-16x^2+576x -------------------------------------------- To complete the square for h(x)=-16x^2+576x, we can find the vertex to help us out We do this by finding the axis of symmetry first -16x^2+576x is in the form ax^2 + bx + c where a = -16 b = 576 c = 0 Plug in a = -16 and b = 576 into the axis of symmetry formula below x = -b/(2a) x = -576/(2*(-16)) x = -576/(-32) x = 18 So the x coordinate of the vertex is 18 Plug this into the function to get the y coordinate of the vertex y = -16x^2+576x y = -16(18)^2+576(18) y = -16(324)+576(18) y = -5184+10368 y = 5184 This is the y coordinate of the vertex. Since the vertex is (18, 5184), we know that (h,k) = (18, 5184) which means h = 18 k = 5184 ------------------------------------------- So you plug a = -16, h = 18 and k = 5184 into the general vertex equation y = a(x-h)^2 + k to get y = a(x-h)^2 + k y = -16(x-18)^2 + 5184

jimthompson5910 (jim_thompson5910):

It's a lot I know, but once you get the hang of it, it's not too bad. In summary, after completing the square y = -16x^2+576x transforms into y = -16(x-18)^2 + 5184

jimthompson5910 (jim_thompson5910):

Let me know when you're ready for the next step

OpenStudy (anonymous):

this is the best response i've gotten so far and now i actually know how to do it! thank you so much! i am ready for the next step! :)

jimthompson5910 (jim_thompson5910):

ok great we want to find the roots of y = -16x^2+576x, which is basically the same thing as saying "solve -16x^2+576x = 0" because y = -16(x-18)^2 + 5184 is equivalent to y = -16x^2+576x, we can solve the equation below -16(x-18)^2 + 5184 = 0 to get the same answers

jimthompson5910 (jim_thompson5910):

So let's solve -16(x-18)^2 + 5184 = 0 for x. We do so by following PEMDAS in reverse to isolate x and get it on its own side \[\Large -16(x-18)^2 + 5184 = 0\] \[\Large -16(x-18)^2 = -5184\] \[\Large (x-18)^2 = \frac{-5184}{-16}\] \[\Large (x-18)^2 = 324\] \[\Large x-18 = \pm\sqrt{324}\] \[\Large x-18 = \sqrt{324} \ \text{ or } \ x-18 = -\sqrt{324}\] \[\Large x-18 = 18 \ \text{ or } \ x-18 = -18\] \[\Large x = 18+18 \ \text{ or } \ x = -18+18\] \[\Large x = 36 \ \text{ or } \ x = 0\]

jimthompson5910 (jim_thompson5910):

So the two solutions of \(\Large -16(x-18)^2 + 5184 = 0\) are... \[\Large x = 36 \ \text{ or } \ x = 0\] That means the two solutions of \(\Large -16x^2+576x = 0\) are... \[\Large x = 36 \ \text{ or } \ x = 0\] since the two equations are equivalent

OpenStudy (anonymous):

THANK YOU SO MUCH but i understand it all besides what the question is really asking for... do you mind if i ask you?

jimthompson5910 (jim_thompson5910):

Sure go ahead

OpenStudy (anonymous):

a projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t)=-16t^2+576t. after how many seconds does the projectile take to reach its maximum height?

jimthompson5910 (jim_thompson5910):

Ok they want the x coordinate of the vertex essentially. In other words, they want the value of t when the projectile is at the max height. Remember how I replaced t with x (I found it easier that way). The x coordinate of the vertex is x = 18. So t = 18 is the time when the projectile is at the max height (ie the vertex) So that's why the answer is t = 18

OpenStudy (anonymous):

THANK YOU SO MUCH @jim_thompson5910

jimthompson5910 (jim_thompson5910):